# VisualPredicator: Learning Abstract World Models with Neuro-Symbolic Predicates for Robot Planning

### v1

@@ C.2.4 Cover Heavy
NSRT-Op1:
Parameters: [?x0:block, ?x1:robot, ?x2:target]
Preconditions: [Holding(?x1:robot, ?x0:block)]
Add Effects: [Covers(?x0:block, ?x2:target), EmptyHand(?x1:robot)]
Delete Effects: [Holding(?x1:robot, ?x0:block)]
Ignore Effects: []
Option Spec: Place(?x0:block, ?x2:target)
NSRT-Op0:
Parameters: [?x0:block, ?x1:robot]
Preconditions: [IsBlack(?x0:block), EmptyHand(?x1:robot)]
Add Effects: [Holding(?x1:robot, ?x0:block)]
Delete Effects: [EmptyHand(?x1:robot)]
Ignore Effects: []
Option Spec: Pick(?x0:block)
@@ C.2.5 Balance
NSRT-Unstack:
Parameters: [?block:block, ?otherblock:block, ?robot:robot]
Preconditions: [Clear(?block:block), DirectlyOn(?block:block, ?otherblock:block), GripperOpen(?robot:robot)]
Add Effects: [Clear(?otherblock:block), Holding(?block:block)]
Delete Effects: [Clear(?block:block), DirectlyOn(?block:block, ?otherblock:block), GripperOpen(?robot:robot)]
Ignore Effects: []
Option Spec: Pick(?robot:robot, ?block:block)
NSRT-Op3:
Parameters: [?block:block, ?otherblock:block, ?robot:robot]
Preconditions: [Clear(?otherblock:block), Holding(?block:block)]
Add Effects: [Clear(?block:block), DirectlyOn(?block:block, ?otherblock:block), GripperOpen(?robot:robot)]
Delete Effects: [Clear(?otherblock:block), Holding(?block:block)]
Ignore Effects: []
Option Spec: Stack(?robot:robot, ?otherblock:block)
NSRT-Op2:
Parameters: [?x0:machine, ?x1:plate, ?x2:plate, ?x3:robot]
Preconditions: [BlocksDistributedEvenly(?x2:plate, ?x1:plate)]
Add Effects: [MachineOn(?x0:machine)]
Delete Effects: []
Ignore Effects: []
Option Spec: TurnMachineOn(?x3:robot, ?x1:plate, ?x2:plate)
NSRT-Op4:
Parameters: [?block:block, ?robot:robot, ?plate:plate]
Preconditions: [ClearPlate(?plate:plate), Holding(?block:block)]
Add Effects: [Clear(?block:block), DirectlyOnPlate(?block:block, ?plate:plate), GripperOpen(?robot:robot)]
Delete Effects: [ClearPlate(?plate:plate), Holding(?block:block)]
Ignore Effects: []
Option Spec: PutOnPlate(?robot:robot, ?plate:plate)
NSRT-PickFromTable:
Parameters: [?block:block, ?robot:robot, ?plate:plate]
Preconditions: [Clear(?block:block), DirectlyOnPlate(?block:block, ?plate:plate), GripperOpen(?robot:robot)]
Add Effects: [Holding(?block:block)]
Delete Effects: [Clear(?block:block), DirectlyOnPlate(?block:block, ?plate:plate), GripperOpen(?robot:robot)]
Ignore Effects: []
Option Spec: Pick(?robot:robot, ?block:block)
@@ C.3 Further Planning Statistics
The average planning node expaneded and wall-time statistics for our approach, alongside other planning approaches, are summarized in the tables.
@@ Appendix D Additional Environment Details
NSRT-Pick:
Parameters: [?block:block]
Preconditions: []
Add Effects: []
Delete Effects: []
Ignore Effects: []
Option Spec: Pick(?block:block)
NSRT-Place:
Parameters: [?block:block, ?target:target]
Preconditions: []
Add Effects: [Covers(?block:block, ?target:target)]
Delete Effects: []
Ignore Effects: []
Option Spec: Place(?block:block, ?target:target)
@@
NSRT-PickFromTable:
Parameters: [?block:block, ?robot:robot]
Preconditions: []
Add Effects: []
Delete Effects: [OnTable(?block:block)]
Ignore Effects: []
@@
[…]
NSRT-PutOnTable:
Parameters: [?block:block, ?robot:robot]
Preconditions: []
Add Effects: [OnTable(?block:block)]
Delete Effects: []
Ignore Effects: []
Option Spec: PutOnTable(?robot:robot)
NSRT-Stack:
Parameters: [?block:block, ?otherblock:block,
?robot:robot]
Preconditions: []
Add Effects: [On(?block:block,
?otherblock:block)]
Delete Effects: []
Ignore Effects: []
Option Spec: Stack(?robot:robot,
?otherblock:block)
NSRT-Unstack:
Parameters: [?block:block, ?otherblock:block,
?robot:robot]
Preconditions: []
Add Effects: []
Delete Effects: [On(?block:block,
?otherblock:block)]
Ignore Effects: []
[…]
It has initial operators:
[…]
NSRT-PickJugFromTable:
Parameters: [?robot:robot, ?jug:jug]
Preconditions: []
Add Effects: []
Delete Effects: []
Ignore Effects: []
Option Spec: PickJug(?robot:robot, ?jug:jug)
NSRT-PlaceJugInMachine:
Parameters: [?robot:robot, ?jug:jug,
?machine:coffee_machine]
Preconditions: []
Add Effects: []
Delete Effects: []
Ignore Effects: []
Option Spec: PlaceJugInMachine(?robot:robot,
?jug:jug, ?machine:coffee_machine)
NSRT-PourFromNowhere:
Parameters: [?robot:robot, ?jug:jug,
?cup:cup]
Preconditions: []
Add Effects: [CupFilled(?cup:cup)]
Delete Effects: []
Ignore Effects: []
Option Spec: Pour(?robot:robot, ?jug:jug,
?cup:cup),
NSRT-TurnMachineOn:
Parameters: [?robot:robot, ?jug:jug,
?machine:coffee_machine]
Preconditions: []
Add Effects: [JugFilled(?jug:jug)]
Delete Effects: []
Ignore Effects: []
Option Spec: TurnMachineOn(?robot:robot,
?machine:coffee_machine),
NSRT-Twist:
Parameters: [?robot:robot, ?jug:jug]
Preconditions: []
Add Effects: []
Delete Effects: []
Ignore Effects: []
Option Spec: Twist(?robot:robot, ?jug:jug)
@@
NSRT-PickFromTable:
Parameters: [?block:block, ?robot:robot,
?plate:plate]
Preconditions: [Clear(?block:block),
DirectlyOnPlate(?block:block, ?plate:plate),
GripperOpen(?robot:robot)]
Add Effects: [Holding(?block:block)]
Delete Effects: [Clear(?block:block),
DirectlyOnPlate(?block:block, ?plate:plate),
GripperOpen(?robot:robot)]
Ignore Effects: []
@@
[…]
NSRT-PutOnPlate:
Parameters: [?block:block, ?robot:robot,
?plate:plate]
Preconditions: [ClearPlate(?plate:plate),
Holding(?block:block)]
Add Effects: [Clear(?block:block),
DirectlyOnPlate(?block:block, ?plate:plate),
GripperOpen(?robot:robot)]
Delete Effects: [ClearPlate(?plate:plate),
Holding(?block:block)]
Ignore Effects: []
Option Spec: PutOnPlate(?robot:robot, ?plate:plate),
NSRT-Stack:
Parameters: [?block:block, ?otherblock:block,
?robot:robot]
Preconditions: [Clear(?otherblock:block),
Holding(?block:block)]
Add Effects: [Clear(?block:block),
DirectlyOn(?block:block, ?otherblock:block),
GripperOpen(?robot:robot)]
Delete Effects: [Clear(?otherblock:block),
Holding(?block:block)]
Ignore Effects: []
Option Spec: Stack(?robot:robot,
?otherblock:block)
NSRT-Unstack:
Parameters: [?block:block, ?otherblock:block,
?robot:robot]
Preconditions: [Clear(?block:block),
DirectlyOn(?block:block, ?otherblock:block),
GripperOpen(?robot:robot)]
Add Effects: [Clear(?otherblock:block),
Holding(?block:block)]
Delete Effects: [Clear(?block:block),
DirectlyOn(?block:block, ?otherblock:block),
GripperOpen(?robot:robot)]
Ignore Effects: []
Option Spec: Pick(?robot:robot,
[…]
NSRT-TurnMachineOn:
Parameters: [?robot:robot, ?machine:machine,
?plate1:plate, ?plate2:plate]
Preconditions: []
Add Effects: [MachineOn(?machine:machine)]
Delete Effects: []
Ignore Effects: []
Option Spec: TurnMachineOn(?robot:robot,
?plate1:plate, ?plate2:plate)

### v2

@@ C.2.4 Cover Heavy
NSRT-Op1: Parameters: [?x0:block, ?x1:robot, ?x2:target] Preconditions: [Holding(?x1:robot, ?x0:block)] Add Effects: [Covers(?x0:block, ?x2:target), EmptyHand(?x1:robot)] Delete Effects: [Holding(?x1:robot, ?x0:block)] Ignore Effects: [] Option Spec: Place(?x0:block, ?x2:target) NSRT-Op0: Parameters: [?x0:block, ?x1:robot] Preconditions: [IsBlack(?x0:block), EmptyHand(?x1:robot)] Add Effects: [Holding(?x1:robot, ?x0:block)] Delete Effects: [EmptyHand(?x1:robot)] Ignore Effects: [] Option Spec: Pick(?x0:block)
@@ C.2.5 Balance
NSRT-Unstack: Parameters: [?block:block, ?otherblock:block, ?robot:robot] Preconditions: [Clear(?block:block), DirectlyOn(?block:block, ?otherblock:block), GripperOpen(?robot:robot)] Add Effects: [Clear(?otherblock:block), Holding(?block:block)] Delete Effects: [Clear(?block:block), DirectlyOn(?block:block, ?otherblock:block), GripperOpen(?robot:robot)] Ignore Effects: [] Option Spec: Pick(?robot:robot, ?block:block) NSRT-Op3: Parameters: [?block:block, ?otherblock:block, ?robot:robot] Preconditions: [Clear(?otherblock:block), Holding(?block:block)] Add Effects: [Clear(?block:block), DirectlyOn(?block:block, ?otherblock:block), GripperOpen(?robot:robot)] Delete Effects: [Clear(?otherblock:block), Holding(?block:block)] Ignore Effects: [] Option Spec: Stack(?robot:robot, ?otherblock:block) NSRT-Op2: Parameters: [?x0:machine, ?x1:plate, ?x2:plate, ?x3:robot] Preconditions: [BlocksDistributedEvenly(?x2:plate, ?x1:plate)] Add Effects: [MachineOn(?x0:machine)] Delete Effects: [] Ignore Effects: [] Option Spec: TurnMachineOn(?x3:robot, ?x1:plate, ?x2:plate) NSRT-Op4: Parameters: [?block:block, ?robot:robot, ?plate:plate] Preconditions: [ClearPlate(?plate:plate), Holding(?block:block)] Add Effects: [Clear(?block:block), DirectlyOnPlate(?block:block, ?plate:plate), GripperOpen(?robot:robot)] Delete Effects: [ClearPlate(?plate:plate), Holding(?block:block)] Ignore Effects: [] Option Spec: PutOnPlate(?robot:robot, ?plate:plate) NSRT-PickFromTable: Parameters: [?block:block, ?robot:robot, ?plate:plate] Preconditions: [Clear(?block:block), DirectlyOnPlate(?block:block, ?plate:plate), GripperOpen(?robot:robot)] Add Effects: [Holding(?block:block)] Delete Effects: [Clear(?block:block), DirectlyOnPlate(?block:block, ?plate:plate), GripperOpen(?robot:robot)] Ignore Effects: [] Option Spec: Pick(?robot:robot, ?block:block)
@@ C.3 Further Planning Statistics
The average planning node expanded and wall-time statistics for our approach, alongside other planning approaches, are summarized in Table 1.
@@ Appendix D Additional Environment Details
NSRT-Pick: Parameters: [?block:block] Preconditions: [] Add Effects: [] Delete Effects: [] Ignore Effects: [] Option Spec: Pick(?block:block) NSRT-Place: Parameters: [?block:block, ?target:target] Preconditions: [] Add Effects: [Covers(?block:block, ?target:target)] Delete Effects: [] Ignore Effects: [] Option Spec: Place(?block:block, ?target:target)
@@
Coffee.
This environment has goal predicates: {CupFilled(?cup:cup)}. We include the predicate JugFilled(?jug:jug) in the initial set of predicates because it was very challenging to have a VLM to determine this especially with the graphics in the simulator. It has initial operators:
NSRT-PickJugFromTable: Parameters: [?robot:robot, ?jug:jug] Preconditions: [] Add Effects: [] Delete Effects: [] Ignore Effects: [] Option Spec: PickJug(?robot:robot, ?jug:jug) NSRT-PlaceJugInMachine: Parameters: [?robot:robot, ?jug:jug, ?machine:coffee_machine] Preconditions: [] Add Effects: [] Delete Effects: [] Ignore Effects: [] Option Spec: PlaceJugInMachine(?robot:robot, ?jug:jug, ?machine:coffee_machine) NSRT-PourFromNowhere: Parameters: [?robot:robot, ?jug:jug, ?cup:cup] Preconditions: [] Add Effects: [CupFilled(?cup:cup)] Delete Effects: [] Ignore Effects: [] Option Spec: Pour(?robot:robot, ?jug:jug, ?cup:cup), NSRT-TurnMachineOn: Parameters: [?robot:robot, ?jug:jug, ?machine:coffee_machine] Preconditions: [] Add Effects: [JugFilled(?jug:jug)] Delete Effects: [] Ignore Effects: [] Option Spec: TurnMachineOn(?robot:robot, ?machine:coffee_machine), NSRT-Twist: Parameters: [?robot:robot, ?jug:jug] Preconditions: [] Add Effects: [] Delete Effects: [] Ignore Effects: [] Option Spec: Twist(?robot:robot, ?jug:jug)